OpenStudy (anonymous):

Write the trigonometric expression in terms of sine and cosine, and then simplify. (1 + tan^2 θ) cos^2 θ

hero (hero):

\[\left(\frac{\cos^2\theta}{\cos^2\theta} + \frac{\sin^2\theta}{\cos^2 \theta}\right)\cos^2\theta\]

hero (hero):

\[\left(\frac{\cos^2\theta + \sin^2\theta}{\cos^2\theta} \right)\cos^2\theta\]

OpenStudy (anonymous):

common denominator right?

hero (hero):

\[\left(\frac{\cos^2\theta + \sin^2\theta}{\cancel{\cos^2\theta}} \right)\cancel{\cos^2\theta}\]

OpenStudy (anonymous):

oic, so will i use the reciprocal identity tan x = sin x/cos x

hero (hero):

\[=\cos^2 \theta + \sin^2 \theta \\=1\]

OpenStudy (anonymous):

neat thanks for your time